OpenStudy (anonymous):

[6] There is a 72% chance that Katie will get married, a 46% chance that she will graduate from college, and a 60% chance that she will marry and graduate. What is the chance that she will get married or graduate?

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

PLEASE HELP!!!!!

OpenStudy (anonymous):

FIRST-write your given values.

OpenStudy (anonymous):

Chance(probability that Katie with get married) = 72% (or 72/100) Chance (probability that Katie will graduate) = 46 % (or 46/100)

OpenStudy (anonymous):

60/100 (probability that Katie will get married and graduate)=60%

OpenStudy (anonymous):

SECOND -->Write your unknown-what you want to solve for.

OpenStudy (anonymous):

i want to solve for the chances she'll either graduate or get married so if 72% is married and 46% is graduate

OpenStudy (anonymous):

Yes-you are solving for the chance/probability Katie will get married OR the chance/probability that she will graduate.

OpenStudy (anonymous):

THIRD-what RULE can you use to solve this problem?

OpenStudy (anonymous):

empirical rule?

OpenStudy (anonymous):

Hint-it is one of the PROBABILITY rules.

OpenStudy (anonymous):

Probability that one OR another event will occur = Probability of desired event #1 occurring + Probability of desired event #2 occurring

OpenStudy (anonymous):

so its adding the married and graduate together

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Yes-correct.

OpenStudy (anonymous):

72+46

OpenStudy (anonymous):

The total probability must never exceed 100%......... YOU MUST USE A SECOND RULE HERE: When two events, #1 and #2, are not mutually exclusive (they can occur at the same time), there is some overlap between these events.........so----------> [The probability that either event #1 OR event #2 will occur] = [Probability of desired event #1 occurring + Probability of desired event #2 occurring] - [THE PROBABILITY THAT BOTH EVENT #1 AND EVENT #2 WILL OCCUR]

OpenStudy (anonymous):

So apply the above rule using your final GIVEN - the one you provided above: probability that Katie will get married and graduate)=60%

OpenStudy (anonymous):

76+42-60

OpenStudy (anonymous):

Yes-awesome!

OpenStudy (anonymous):

58

OpenStudy (anonymous):

Correct-great job!

OpenStudy (anonymous):

thank you!!!!!!!!

OpenStudy (anonymous):

could you help with another one??

OpenStudy (anonymous):

me*

OpenStudy (anonymous):

hey can you help with more question please

OpenStudy (anonymous):

[7] Find the radian measure for 2100 and its associated coordinate point on the unit circle. Then find the sin, cos, and tangent. 210(π/180) = 7 π/6 radians Notice the 210 degrees brings ( - v3/2, -1/2) The Sin=y/rad = -1/2 / 1 = -1/2 = The Cos=x/rad = -v3/2 / 1 = The Tan=y/x = -1/2 /-v3/2 = -1/ v3 =

OpenStudy (anonymous):

I am not sure how to do this one-sorry!